zunairah:

At the top of a 75 m hill is a car weighing 91 N. What is its potential energy? The same car (mass of 9.3 kg) goes down the hill and at the bottom is traveling 37m/s. What is its kinetic energy ?

Florisalreadytaken:

for the first one, we have to find potential energy -- we can simply do that by using the formula \[ \text{potential energy} (U) = mgh\]

Florisalreadytaken:

the formula we have to use for the 2nd question is: \( \text{kinetic energy}= \frac{1}{2} m v^2 \)

Florisalreadytaken:

and we are given \[ m=9.3kg\] \[ v=37 m/s \]

Florisalreadytaken:

just substitute the numbers, and there you go! :D

zunairah:

for the first one I got 655473 for PE

zunairah:

and for the 2nd one i got 12731.7

zunairah:

I think its right

Florisalreadytaken:

they're both wrong...

zunairah:

ummm

Florisalreadytaken:

are you sure the numbers you inserted in the calculator were right?

zunairah:

yeah..

zunairah:

what was i supposed to get

zunairah:

66885 for the first one?

zunairah:

344.1 for the second?

Florisalreadytaken:

firstly, turn newtons into kg

zunairah:

oh

zunairah:

idk how to do this

zunairah:

hOw?

zunairah:

what even is this

Florisalreadytaken:

lol am jk

Florisalreadytaken:

so \[m=9.3 kg\] \[g=9.8m/s^2\] \[h=75m\] thus \[p.e = 6835.5 J\]

Florisalreadytaken:

as for the 2nd one \[ \text{kinetic energy}= \frac{1}{2} 9.3kg \times 37m/s \; ^2 \]

Florisalreadytaken:

\[ \text{kinetic energy} = \color{lightskyblue}{6365.85} J\]

zunairah:

thx